OpenStudy (anonymous):

Factor completely: 18x2 − 21x − 15

OpenStudy (aroub):

3( 2x+1)(3x-5)

OpenStudy (anonymous):

\[(6x - 3) (3x +5)\]

myininaya (myininaya):

\[3(6x^2-7x-5)=3(6x^2-10x+3x-5) \] \[=3(2x(3x-5)+1(3x-5))=3(3x-5)(2x+1)\] aroub's answer is right but bob has his signs wrong but nice attemp or we can say \[(6x+3)(3x-5)\]

OpenStudy (anonymous):

Woops! Nice spot myininaya :)

myininaya (myininaya):

:) thanks